OpenStudy (anonymous):

There are two factors of -36 such that one factor is 11 less than half of the other factor. Choose all the pairs of these factors. -2 and 18 -6 and 6 3 and -12 4 and -9

OpenStudy (jdoe0001):

well, we know that the set is (a , b), and that say "a" is 11 less than half of "b" or \(a = \cfrac{b}{2}-11\)

OpenStudy (jdoe0001):

so, try out the sets, plug the values for "b", see if you get a matching "a"

OpenStudy (jdoe0001):

let's try the 1st one -2 and 18 if b = 18, then a = -2 let's see \(\large a = \frac{18}{2}-11\) what do you get for "a"?

OpenStudy (phi):

or solve b^2/2 - 11b = -36 for b a * b= -36 and a= b/2 -11

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

is that correct?

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

so y = 18, gave x = -2 so, that's the correct set :)